OpenStudy (anonymous):

What is the interval of convergence for this power series: the sum from n=1 to infinity of ((x-5)^n)/((n!)5^n)?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{(x-5)^{n}}{ (n!)5^{n} }\]

OpenStudy (anonymous):

The answer is \[( -\infty , \infty ) \] I just need to know how to get there.

terenzreignz (terenzreignz):

Let's try using the Ratio Test ^_^

terenzreignz (terenzreignz):

\[\Large \lim_{n\rightarrow \infty}\left[\frac{(x-5)^{n+1}}{(n+1)!5^{n+1}}\cdot \frac{n!5^n}{(x-5)^n}\right]\]

terenzreignz (terenzreignz):

\[\Large = \lim_{n\rightarrow \infty}\left[\frac{x-5}{5(n+1)}\right]\] Unless I'm mistaken with the cancellations :D

OpenStudy (anonymous):

That's what I got so far

terenzreignz (terenzreignz):

Hang on, let's try putting more detail into this... \[\Large \lim_{n\rightarrow \infty}\left[\frac{\color{blue}{(x-5)^{n}}(x-5)}{(n+1)\cdot \color{red}{n!}\cdot\color{green}{5^{n}}\cdot5}\cdot \frac{\color{red}{n!}\color{green}{5^n}}{\color{blue}{(x-5)^n}}\right]\] So, the blues, greens, and reds cancel out, leaving \[\Large = \lim_{n\rightarrow \infty}\left[\frac{x-5}{5(n+1)}\right]\]

terenzreignz (terenzreignz):

This limit goes to zero as n goes to infinity... did you notice? Don't forget that x is taken as a constant as far as n is concerned ^_^ \[\Large = \lim_{n\rightarrow \infty}\left[\frac{x-5}{5(n+1)}\right]=\color{red}0\]

terenzreignz (terenzreignz):

Which means, by the ratio test, this converges (absolutely) regardless of the value of x :D

terenzreignz (terenzreignz):

(Regardless of the value of x... because the limit does not involve x )

OpenStudy (anonymous):

So if we do the ratio test and it is less than 1, then don't solve the inequality?

terenzreignz (terenzreignz):

If we do the ratio test and it is less than one, then you know that the series converges regardless of x. If on the other hand, the limit involves x somehow, that's when you solve an inequality. for instance, say, if the limit was \[\frac{x+1}3\] then you'd have to solve \[\Large \left|\frac{x+1}{3}\right|<1\] to get the radius and interval of convergence :3

OpenStudy (anonymous):

Typically you're solving the limit as n approaches infinity, so isn't x usually pulled out of the limit?

OpenStudy (anonymous):

Anyways, I appreciate your help!

terenzreignz (terenzreignz):

Sorry, must have dozed off. and just treat x as a constant and work from there :)